OpenStudy (anonymous):

How do I put this into slope intercept form? (2, 120) (4, 240) 120-240/2-4 =60

OpenStudy (campbell_st):

well thats the slope m the point slope form of a line is y = mx + b m = slope and b = y-intercept to calculate b, choice a point say (2, 120) and substitute into y = 60x + b to find the value of b when you have done that you'll have your line in point slope form.

OpenStudy (anonymous):

I did that and this is what I get: y=mx+b 120=60(2)+b 120=120+b 0=b ?? I'm just confused.

OpenStudy (campbell_st):

thats correct so whats th equation of the line..?

OpenStudy (campbell_st):

remember its y = mx + b you've found m = 60 and b = 0 just substitute for the equation

OpenStudy (anonymous):

??? I don't know... so... y = 60x + 0 ????

OpenStudy (anonymous):

or is it just y=60x ??? Thank you for your help, by the way, :)

OpenStudy (campbell_st):

thats it.... to check its correct use the other point (4, 240) subsitiute x = 4 and you should get y = 240